2x^2+10x+12=0 Can someone please explain this to me with steps. I don't understand it quite yet, steps and an explanation would be very helpful. I'm very, very new to this.

Factor
Answer #1

Solution:

Comparing 2x2 + 10x + 12 0    with \small ax^{2}+bx+c=0    we get a= 2, b= 10, c=12

Now, using the quadratic formula,

\small x=\frac{-b\pm \sqrt{b^{2}-4ac}}{2a}

-10± V102-4 x 2 x 12

-10± v 100-96

10± V4

-10± 2

_10+ 2 一10-2

\small \therefore x=\frac{-8}{4}\: \: \: \: or\: \: \: x=\frac{-12}{4}

\small \therefore x=-2\: \: \: \: or\: \: \: x=-3